OpenStudy (anonymous):

Simplify the rational expression. 4x^2 − 20x/x^3 − 125

OpenStudy (earthcitizen):

\[x/(x ^{2}+5x+25)\]

OpenStudy (anonymous):

lol, wrong

OpenStudy (earthcitizen):

lol, how ?

OpenStudy (earthcitizen):

\[(4x^2 − 20x)/(x^3 − 125) \] divide num by 4, \[x(x-5)/(x^3 - 5^3) = x(x-5)/((x-5)(x^2+5x+25) = x/(x^2+5x+25)\]